A particle is projected horizontally with a speed 20m/s from the top of a tower.after what time will the velocity of the particle be 45 degree angle from the initial direction of projection
1
Expert's answer
2021-09-13T11:02:09-0400

"\vec v_x= 20\frac{m}{s}"

"g = 9.8 \frac{m}{s^2}"

"\vec v_y =\frac{gt^2}{2}"

"\vec{v} =\vec v_x+\vec v_y"

"\alpha= \angle(\vec v,\vec v_x)=45 \degree"

"\vec v_y =\tan \alpha*\vec{v_x}= v_x=20\frac{m}{s^2}"

"t = \sqrt{\frac{2v_y}{g}}=\sqrt{\frac{2*20}{9.8}}=2.02s"

"\text{Answer: }2.02s"